OpenStudy (anonymous):

the formula for degrees celsius is c=5/9 (f-320 where f stands for degrees fahrenheit part 1: solve the equation for f show all steps part2: determine how many degrees fahrenheit.

OpenStudy (anonymous):

i honestly don't know how to do this at all.

OpenStudy (anonymous):

\[C=\frac{5}{9}(F-32)\] solve for F in two steps

OpenStudy (anonymous):

1) multiply both sides by \(\frac{9}{5}\) to get \[\frac{9}{5}C=F-32\]

OpenStudy (anonymous):

2) add \(32\) to get \[\frac{9}{5}C+32=F\]

OpenStudy (anonymous):

that is done

OpenStudy (anonymous):

wow that was simple my friends were saying distribute.but that didn't sound right. ha thank you very much!

OpenStudy (anonymous):

you could that would be way way more work

OpenStudy (anonymous):

yw